The line x=y=z-1 intersects the cone x^2+y^2+z^2+2yz+2zx+2xy=0 at exactly one point
1
Expert's answer
2020-11-02T18:51:18-0500

Given equation of line is

Let

"x=y=z-1=t"


we get

"x=t\\ y=t\\ z=1+t"


Given Equation of cone is

"x^2+y^2+z^2+2xy+2yz+2xz=0\\\to t^2+t^2+(1+t)^2+2t^2+2t(1+t)+2t(1+t)=0\\\to 9t^2+6t+1=0"


"\to (3t+1)^2=0\\\to t=-\dfrac{1}{3}"


so we get,

"x=-\dfrac{1}{3}\\ y=-\dfrac{1}{3}\\z=1-\dfrac{1}{3}=\dfrac{2}{3}"



So given line intersect with cone at "(\dfrac{-1}{3},\dfrac{-1}{3},\dfrac{2}{3})" .